Low Exposure

Socioeconomic Status (SES)

Disease Status

Low SES

High SES

Has AD

52

86

No AD

11,100

4,042

High Exposure

Socioeconomic Status (SES)

Disease Status

Low SES

High SES

Has AD

10

157

No AD

4,081

6,321

Individuals who are low SES are [x] times more likely to have an AD than individuals who have higher SES

*Do not round up for your calculations and use X.X format
Answer #1

Individuals who are SES with AD = 52 + 10 = 62

Individuals who are high SES with AD = 86 + 157 = 243

So , 62÷ 243 = 0.255

Therefore individuals who are low SES are 0.25 times more likely to have an AD than individuals who have higher SES .